OpenStudy (anonymous):

Solve. Do checks. Explain. y = 2x - 3 -y = x Systems Of Elimination/Addition (I'm a visual learner)

geerky42 (geerky42):

-y = x so y = -x Replace y to -x in first equation. Now solve the first equation.

geerky42 (geerky42):

Not visual enough?

OpenStudy (mathstudent55):

To solve a system of equations by substitution, solve one equation for one variable, and substitute in the other. Then solve the equation for one variable. Once you have that variable, substitute its value into one of the original equations to get the other variable. Since the second equation is solved for x, x = -y, where you see x in the first equiation, insert -y. y = 2x - 3 We know x = -y, so y = 2(-y) - 3 y = -2y - 3 3y = -3 y = -1 Now substitute -1 in for y in the first equation: y = 2x - 3 -1 = 2x - 3 2 = 2x 1 = x Solution: x = 1, y = -1

OpenStudy (anonymous):

^ Solve systems by ADDITION...

geerky42 (geerky42):